OpenStudy (anonymous):

what is y= -7+6x in standard form?

OpenStudy (anonymous):

-6x+y=7 AX+BY=C

OpenStudy (anonymous):

Standard form Ax+By=C y=-7+6x A should be positive. I'd write 6x-y=-7

OpenStudy (anonymous):

A doesn't have to be positive, it just stands for an integer.

OpenStudy (anonymous):

Otherwise B would be positive too and you couldn't get a correct answer.